# Measuring AI R&D Automation

### 04 Mar 2026 v1

@@ 1 Introduction
The effects of AI R&D automation (AIRDA) could be significant, but might cut in different directions (Figure˜1). AIRDA could accelerate AI progress, bringing forward AI’s benefits but also hastening the arrival of destructive capabilities, including those related to weapons of mass destruction, or other forms of disruption such as unemployment (Bengio et al., 2026). Whether this acceleration would be overall beneficial remains unclear. Key uncertainties include whether AIRDA would accelerate defensive capabilities more than offensive ones, whether safety research will keep pace with capabilities research, and whether the human institutions can adapt to the accelerated pace of progress (Bernardi et al., 2025; Vaintrob and Cotton-Barratt, 2025; Kembery and Ammann, 2025).
@@
Existing data on AIRDA has significant gaps. First, many uncertainties remain about the extent of AIRDA. Rapidly improving benchmark results indicate at least some progress (Jimenez et al., 2023; Chan et al., 2025; Starace et al., 2025; Wijk et al., 2025; Anthropic, 2026b), [^2] At the time of writing, the most advanced model had a 80% success rate on tasks that take human expert coders 1 hour and 10 minutes to complete (Metr, 2025b). but it is unclear how directly such results translate to productivity boosts given real-world integration frictions (Becker et al., 2025; Dell’Acqua et al., 2023; Brynjolfsson et al., 2023; Noy and Zhang, 2023; Narayanan and Kapoor, 2025; Becker et al., 2026). Current benchmarks also mainly focus on software engineering rather than other parts of the research pipeline, such as generation and prioritization of research ideas. Second, there is little data about the consequences of AIRDA, rather than simply its extent. For example, it is unclear how productivity improvements differ between safety teams and other teams, or how often AI-introduced errors escape or could escape human review.
@@
Companies Track differential progress between safety and capabilities research Metric #1 (AI R&D evaluations) and Metric #6 (staff surveys) are most straightforward. Metric #2 (human–AI comparison RCTs) can also track differential progress, but requires significantly more effort.
@@
Third parties Create tooling and design surveys Metric #8 (researcher time allocation) tooling, and surveys for Metric #6 (self-reported productivity boosts) and Metric #7 (AI use in high-stakes decisions).
@@ 2 What is AI R&D?
AI R&D encompasses the activities involved in developing and improving AI systems, including both capabilities research and safety and security research. We explain activities involved in the current AI R&D process (roughly following Owen (2024)) and comment on how AI could plausibly automate its various stages.
@@
Running experiments. Running experiments can sometimes be as simple as running the code. Other times, especially for large-scale experiments, AI researchers have to efficiently distribute experiments across multiple machines and monitor experiments for signs of failure. Monitoring large-scale training runs can be particularly tricky, involving heuristics and on-the-fly interventions (see Marin (2025) for an example). AI systems can potentially assist with monitoring and troubleshooting training runs, but data on effectiveness remains limited.
@@ 3 Potential Implications of AI R&D Automation
To identify informative metrics for measuring AIRDA, we first discuss its potential implications for AI progress and oversight. For AI progress, AIRDA could affect when certain kinds of capabilities emerge and how we respond to them (Figure˜2). For oversight, AIRDA could either widen or narrow the gap between how much oversight of AI R&D is needed and how much is achieved (Figure˜3).
@@ 3.1 AI Progress
  • Safety and security research could facilitate the use of defensive capabilities (e.g., by ensuring that systems follow user instructions) or directly constrain offensive capabilities (e.g., through misuse safeguards). However, safety may be harder to accelerate than capabilities because safety is more difficult to measure. For example, attempts from AI systems to hide misbehaviour could make it harder to assess whether safety is actually improving (Greenblatt et al., 2024; Meinke et al., 2025). If so, accelerated capabilities progress could outstrip the development of necessary safeguards, which could result in a net harm.
@@ 4.2.1 Metric #6: Staff views on AI use and productivity boosts
Feasibility: High. Implementation requires only periodic surveys with straightforward questions. Most organizations already have infrastructure for such data collection. As further evidence of feasibility, the Claude Opus 4.5 System Card (Anthropic, 2025b) includes data from an internal survey that asks whether Opus “could completely automate a junior ML researcher.” The main challenge is maintaining response rates and avoiding survey fatigue.
@@ 4.4.1 Metric #11: Headcount of AI researchers and the distribution of their performance
  • Non-linearity: We should expect human labour to become more valuable before full automation because AI systems can augment human labour (Gans and Goldfarb, 2026). As such, headcount could remain high or even increase as long as human-dependent activities bottleneck full automation (e.g., providing high-level research direction) (Toner et al., 2026).
@@ 4.4.2 Metric #12: Distribution of compute usage
Feasibility: Moderate to high. Compute is likely to be meticulously tracked at frontier AI companies for cost management and infrastructure planning. For example, pre-training workloads require months-long planning and dedicated hardware reservations. The main implementation challenge is categorizing uses of compute. Public estimates based on dollar spend could be possible (You, 2025), but would not track fine-grained details about how compute is used.
@@ 4.4.3 Metric #13: Capital share of AI R&D spending
Description: This metric measures compute expenditure as a percentage of total AI R&D spending. [^6] A broader variant of this metric is the network-adjusted capital share Trammell (2026), which would measure the extent to which the economy as a whole could sustain itself without human labour. It is calculated as:

### 05 Mar 2026 v2

@@ 1 Introduction
The effects of AI R&D automation (AIRDA) could be significant, but might cut in different directions (Figure˜1). AIRDA could accelerate AI progress, bringing forward AI’s benefits but also hastening the arrival of destructive capabilities—including those related to weapons of mass destruction—or other forms of disruption such as unemployment (Bengio et al., 2026). Whether this acceleration would be overall beneficial remains unclear. Key uncertainties include whether AIRDA would accelerate defensive capabilities more than offensive ones, whether safety research will keep pace with capabilities research, and whether the human institutions can adapt to the accelerated pace of progress (Bernardi et al., 2025; Vaintrob and Cotton-Barratt, 2025; Kembery and Ammann, 2025).
@@
Existing data on AIRDA has significant gaps. First, many uncertainties remain about the extent of AIRDA. Rapidly improving benchmark results indicate at least some progress (Jimenez et al., 2023; Chan et al., 2025; Starace et al., 2025; Wijk et al., 2025; Anthropic, 2026b), [^2] At the time of writing, the most advanced model had a 80% success rate on tasks that take human expert coders 1 hour and 10 minutes to complete (Metr, 2025b). but it is unclear how directly such results translate to productivity boosts given real-world integration frictions (Becker et al., 2025; Dell’Acqua et al., 2023; Brynjolfsson et al., 2023; Noy and Zhang, 2023; Narayanan and Kapoor, 2025; Becker et al., 2026). Current benchmarks also mainly focus on software engineering rather than other parts of the research pipeline, such as generation and prioritization of research ideas. Second, there is little data about the consequences of AIRDA, rather than simply its extent. For example, it is unclear whether productivity improvements differ between safety teams and other teams, or how often AI-introduced errors escape human review.
@@
Companies Track differential progress in automating safety vs. capabilities research Metric #1 (AI R&D evaluations) and Metric #6 (staff surveys) are most straightforward. Metric #2 (human–AI comparison RCTs) can also track differential progress, but requires significantly more effort.
@@
Third parties Create tooling and design surveys Time-tracking software for Metric #8 (researcher time allocation), and surveys for Metric #6 (self-reported productivity boosts) and Metric #7 (AI use in high-stakes decisions).
@@ 2 What is AI R&D?
AI R&D encompasses the activities involved in developing and improving AI systems, including both capabilities research and safety and security research. Roughly following Owen (2024), we explain activities involved in the current AI R&D process and comment on how AI could plausibly automate its various stages.
@@
Running experiments. Running experiments can sometimes be as simple as running the code. Other times, especially for large-scale experiments, AI researchers have to efficiently distribute experiments across multiple machines and monitor experiments for signs of failure. Monitoring large-scale training runs can be particularly tricky, involving heuristics and on-the-fly interventions (Marin, 2025). AI systems can potentially assist with monitoring and troubleshooting training runs, but data on effectiveness remains limited.
@@ 3 Potential Implications of AI R&D Automation
To identify informative metrics for measuring AIRDA, we first discuss its potential implications for AI progress and oversight. For AI progress, AIRDA could affect when certain kinds of capabilities emerge and how we respond to them (Figure˜2). For oversight, AIRDA could affect the gap between how much oversight of AI R&D is needed and how much is achieved (Figure˜3).
@@ 3.1 AI Progress
  • Safety and security research could facilitate the use of defensive capabilities (e.g., by ensuring that systems follow user instructions) or directly constrain offensive capabilities (e.g., through misuse safeguards). However, safety may be harder to accelerate than capabilities because safety is more difficult to measure. For example, attempts from AI systems to hide misbehaviour could make it hard to assess whether safety is actually improving (Greenblatt et al., 2024; Meinke et al., 2025). If so, accelerated capabilities progress could outstrip the development of necessary safeguards, which could result in a net harm.
@@ 4.2.1 Metric #6: Staff views on AI use and productivity boosts
Feasibility: High. Implementation requires only periodic surveys with straightforward questions. Most organizations already have infrastructure for such data collection. As further evidence of feasibility, the Claude Opus 4.5 System Card includes data from an internal survey that asks whether Opus “could completely automate a junior ML researcher” (Anthropic, 2025b). The main challenge is maintaining response rates and avoiding survey fatigue.
@@ 4.4.1 Metric #11: Headcount of AI researchers and the distribution of their performance
  • Non-linearity: We should expect human labour to become more valuable before full automation because AI systems can augment human labour (Gans and Goldfarb, 2026). As a result, headcount could remain high or even increase as long as human-dependent activities bottleneck full automation, such as providing high-level research direction (Toner et al., 2026).
@@ 4.4.2 Metric #12: Distribution of compute usage
Feasibility: Moderate to high. Compute is likely to be meticulously tracked at frontier AI companies for cost management and infrastructure planning. For example, pre-training workloads require months-long planning and dedicated hardware reservations. The main implementation challenge is categorizing uses of compute. Public estimates based on dollar spend are possible (You, 2025), but would not track fine-grained details about how compute is used.
@@ 4.4.3 Metric #13: Capital share of AI R&D spending
Description: This metric measures compute expenditure as a percentage of total AI R&D spending. [^6] A broader variant of this metric is the network-adjusted capital share (Trammell, 2026), which would measure the extent to which the economy as a whole could sustain itself without human labour. It is calculated as: